OpenStudy (anonymous):

let x1, x2,....xn be a random sample from a population that follows the following pdf

OpenStudy (anonymous):

f(x)= 0 for x<1 (Theta)x^-(theta+1) for x greater than or equal to 1

OpenStudy (anonymous):

\[(\theta \hat)= 1/n \sum_{i=1}^{n} \ln X1 \]

OpenStudy (anonymous):

is an unbiased estimator for theta